OpenStudy (anonymous):

how to find the derivative using the quotient rule of (s)/(s^1/2-1)

OpenStudy (anonymous):

HELLLLLLLLP

OpenStudy (anonymous):

Wow there are two divisions?

OpenStudy (anonymous):

Oh wait, nevermind, I see.

OpenStudy (anonymous):

Okay so what is derivative of \(s\)?

OpenStudy (anonymous):

what is the derivative of the whole equation.

OpenStudy (anonymous):

What is the derivative of \(\sqrt{s}-1\)

OpenStudy (anonymous):

You have to find the derivative of the parts first.

OpenStudy (anonymous):

\[ \left( \frac fg\right)'=\frac{f'g-fg'}{g^2} \]

OpenStudy (anonymous):

This is quotient rule. Right now \(f=s\) and \(g=\sqrt{s}-1\). We need \(f'\) and \(g'\).